Solve by using the strategies for solving application problems.
A painter can paint a kitchen in 15 hours. An apprentice can paint the same kitchen in 30 hours. If they worked together, how long would it take them to paint the kitchen?
=h

Answer by rothauserc(4718)

working together
1/15 + 1/30 = 1/h
2/30 + 1/30 = 1/h
3/30 = 1/h
1/10 = 1/h
h = 10 hours
working together it will take them 10 hours to complete the job